OpenStudy (anonymous):

Find the value of y so that the line passing through (9, 0) and (3, y) has a slope of -1/2.

OpenStudy (johnweldon1993):

Well you know slope is \[\large m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] Your 2 points are given and we know slope is -1/2 9 , 0 = x1 , y1 3 , y = x2 , y2 So plugging all this in we have \[\large -\frac{1}{2} = \frac{ y - 0 }{ 3 - 9 }\] Now just solve for 'y'